# Occam’s LGS: A simple approach for Language Gaussian Splatting

### v1

@@ 6 Evaluation
We observe that different object queries yield varying ranges of relevancy scores, which makes it challenging to set a consistent threshold for object segmentation across query types, as shown in Figure 7. To address this limitation, we propose a dynamic thresholding approach. In other words, each query has a different threshold value. For each query, we evaluate different threshold values by incrementing with a step size of 0.01. At each threshold, we create a binary mask by selecting regions where the relevancy score exceeds the current threshold value, and compute the average relevancy score within this masked area. These average relevancy scores at different thresholds are plotted in Figure 9, shown by the increasing trend lines.
Our goal is to identify a "stable region" where variations in the threshold minimally affect both the mask size and average relevancy score. Specifically, we find the region where the absolute gradient of the average relevancy score (the rate of change of average relevancy score with respect to different thresholds) is below the chosen stability threshold, as shown in Figure 8. We then select the middle value from this stable range as our final threshold for each different query, as illustrated in Figure 9. The presence of a stable region indicates an object with consistent relevancy scores across threshold changes. In cases where multiple stable regions exist, we select the rightmost one as it typically represents the highest confidence predictions. We apply this improved evaluation methodology to the 3D-OVS dataset to ensure fair comparison across different query types.
Figure 8: Visualization of the absolute gradient of average relevancy scores. The gradient measures the rate of change in average relevancy scores across different thresholds. The green dotted line indicates the stability threshold, with lower gradient values corresponding to more stable regions. This plot corresponds to the "weaving basket" query in room scene.
Figure 9: Relationship between score thresholds, normalized mask sizes, and relevancy scores for the "weaving basket" query in room scene. The plot shows how mask sizes decrease (downward trends with dotted lines) and average relevancy scores change (upward trends with smooth lines) as the threshold increases. Different hierarchical levels are represented by distinct colors, with the shaded region indicating the optimal stable region identified by our stability threshold.
7 More Visualizations
We show more examples on the 3D-OVS dataset in Figure 1011, and  12.

### v2

@@ 8 Evaluation on 3DOVS Dataset
We observe that different object queries yield varying ranges of relevancy scores, which makes it challenging to set a consistent threshold for object segmentation across query types, as shown in Figure 8. To address this limitation, we propose a dynamic thresholding approach. For each query, we evaluate a range of threshold values with a step size of 0.01. At each threshold, we create a binary mask by selecting regions where the relevancy score exceeds the current threshold value, and compute the average relevancy score within this masked area. These average relevancy scores at different thresholds are plotted in Figure 10, shown by the increasing trend lines.
Our goal is to identify a ”stable region” where variations in the threshold minimally affect both the mask size and average relevancy score. Specifically, we find the region where the absolute gradient of the average relevancy score (the rate of change of average relevancy score with respect to different thresholds) is below the chosen stability threshold, as shown in Figure 9. We then select the middle value from this stable range as our final threshold for each different query, as illustrated in Figure 10. The presence of a stable region indicates an object with consistent relevancy scores across threshold changes. In cases where multiple stable regions exist, we select the highest confidence prediction. We apply this improved evaluation methodology to the 3D-OVS dataset to ensure fair comparison across different query types.
Figure 9: Visualization of the absolute gradient of average relevancy scores. The gradient measures the rate of change in average relevancy scores across different thresholds. The green dotted line indicates the stability threshold, with lower gradient values corresponding to more stable regions. This plot corresponds to the ”weaving basket” query in the room scene.
Figure 10: Relationship between score thresholds, normalized mask sizes, and relevancy scores for the ”weaving basket” query in room scene. The plot shows how mask sizes decrease (downward trends with dotted lines) and average relevancy scores change (upward trends with smooth lines) as the threshold increases. Different hierarchical levels are represented by distinct colors, with the shaded region indicating the optimal stable region identified by our stability threshold.
9 Additional Visualizations
9.1 LERF Visualization
We show additional visualization of the semantic features on LERF Dataset in Figure 11.
Figure 11: PCA visualization of semantic features on LERF Dataset. (a) Original RGB rendering. (b) PCA visualization of our rendered semantic feature maps. (c) Initial semantic feature maps extracted using CLIP from SAM segmentations.
9.2 3DOVS Visualization
We show more examples on the 3D-OVS dataset in Figure 1213, and  14.